Enchanced Binding in non-relativistic QED

Semjon Vugalter University of Munich

For a spinless particle coupled to a photon field we prove that even if the Schroedinger operator without photon field does not have eigenvalues the system can have a ground state. We describe the coupling by means of the Pauli-Fierz Hamiltonian and our result holds in the case where the coupling constant is small.